OpenStudy (stormoshawty):

write y = ( -3/4)x + 3 in standard form.

OpenStudy (stormoshawty):

That doesn't make any sense to me. @shamil98 work me through this.

hartnn (hartnn):

do you know what standard form of a line is ?

OpenStudy (stormoshawty):

Yep.

hartnn (hartnn):

good, so to eliminate the deniminator, start by multiplying 4 on both sides

hartnn (hartnn):

**denominator

hartnn (hartnn):

\(y \times 4 = (-3/4)x \times4 +3 \times 4 \\....\)

OpenStudy (nincompoop):

still needing help, @Stormoshawty ?

OpenStudy (nincompoop):

Recall that the slope-intercept form of a line is: y = mx + b. To change this into standard form, we start by moving the x-term to the left side of the equation. This is done by subtracting mx from both sides. We now have the equation, -mx + y = b. The coefficient of the x-term should be a positive integer value, so we multiply the entire equation by an integer value that will make the coefficient positive, as well as, all of the coefficeints integers. This gives us the standard form: Ax + By = C. http://jwilson.coe.uga.edu/emt668/emat6680.2002/jackson/chapter%205%20lesson%20plan/day6.html

OpenStudy (shamil98):

Also you want you want whole numbers, which is why hartnn said to multiply by 4 to both sides.

OpenStudy (stormoshawty):

You're crazy if you think i'm reading all that.